# Assessment of DKIST/VTF Capabilities for the Detection of Local Acoustic Source Wavefronts

### v1

@@ 6 Observational Strategies – Implications for DKIST/VTF
The two measures, $\lambda^{*}(t)$ and $\lambda^{**}$ motivate different observing strategies. The first is an estimate for the best wavelength sequence to employ to follow the wavefront propagation, while the second is an estimate of the best overall choice of wavelength at which to perform monochromatic imaging at high cadence. Observing at $\lambda^{**}$ is roughly equivalent to observing the wavefronts as they propagate horizontally at the height of maximum sensitivity. Observing sequentially over the wavelengths $\lambda^{*}(t)$, on the other hand, allows collection of a cube of data that can be post-processed to find the signatures of wavefronts as they propagate upward with height, measuring, in order, at the most sensitive wavelength at each height. For a more robust measurement, the individual $\lambda^{*}$ maps in multi-wavelength observations could be shifted in time, stacked, and summed to obtain the mean signal over all heights. In this operation, the wave signal adds coherently, while the variation in granulation with height contributes to some cancellation of the background and consequent enhancement of the propagating wavefront signal-to-noise.
@@
The VTF provides full-field imaging spectroscopy across 520087005200–8700  Å, with a $60^{\prime\prime}\times 60^{\prime\prime}$ field of view sampled at $0.014^{\prime\prime}$ per pixel, corresponding to about $10.5$ km on the solar surface. This sampling is comparable to the 16 km horizontal grid spacing of the MURaM simulation analyzed for this paper, and allows access to the spatial scales over which acoustic wavefronts remain coherent and may be recoverable by temporal differencing (Bahauddin and Rast, 2023). The spectral resolving power of the VTF is $R\sim 10^{5}$ (e.g., $\sim$0.06 Å resolution with $\sim$0.03 Å sampling near 6000 Å) over a free spectral range of $\sim$10 Å. In detail, the VTF spectral sampling width is about 0.023 Å at 5250 Å and 0.033 Å at 6302 Å (Tritschler et al., 2016; Schmidt et al., 2016). These band passes are over-plotted in the insets of Figure 6.
@@
As discussed above, the distributions of $\lambda^{**}$, as determined from the individual atmospheres of the annulus of interest, have widths of $\sigma_{\lambda^{**}}=0.0031$ Å, $\sigma_{\lambda^{**}}=0.0028$ Å, and $\sigma_{\lambda^{**}}=0.0046$ Å  for Fe I 5250.2, Fe I 5250.6, and Fe I 6302.5, respectively. Thus, the range of $\lambda^{**}$ values found for each of these lines lies well within one spectral sampling element of the VTF. VTF observations centered on $\lambda^{**}$ or $\lambda^{*}$ will capture a relatively broad range of wavelengths and consequently sample a range of atmospheric heights. From Figure 6, it is clear that, in the case of Fe I 5250.2 for example, centering the VTF passband on $\lambda^{\ast\ast}$ will include significant contributions from the continuum. Those contributions are dominated by granulation noise that will likely overpower the wave signal in monochromatic observations. More generally, the range of atmospheric heights encompassed by the VTF bandpass may smear the wave signal. However, the blue wing of Fe I 5250.6, and even more so that of Fe I 6302.5, are shallower and thus less contaminated by contributions away from the wavelength of maximum sensitivity (see insets of Figure 6). Here we focus our observational plans on the Fe I 6302.5 line, though careful convolution with both VTF spectral bandpass and the DKIST point spread function (PSF) is required to fully assess the potential of each line for local wavefront discovery.
@@ 6.1 Monochromatic Observations
To comply with the requirements of the temporal difference filter, the cadence of the monochromatic observations should be at or below the critical sampling rate $\Delta t=\Delta x/v_{\mathrm{ph}}$, where $\Delta x$ is the spatial sampling width and $v_{\mathrm{ph}}$ is the phase speed of the wavefront. The phase speed of the wavefront at z=0.152z=0.152\,Mm, the midpoint of the Fe I 6302.5 $\lambda^{**}$ height, is $10.5$ km/s. Critical sampling of the wavefront thus occurs at a cadence of 1 s. The practical viability of this cadence is governed by the signal-to-noise requirements based on the intensity fluctuations caused by acoustic perturbations.
@@
Quantitative guidance on the achievable cadence vs. signal-to-noise trade space is assessed by the VTF Instrument Performance Calculator (IPC, National Solar Obseervatory, 2026). For monochromatic (single-wavelength) intensity imaging, the IPC indicates that an SNR of 235 can be achieved at a cadence of 0.04 s, while relaxing the cadence to 1.04 s increases the SNR to 1200. While our analysis is for a single representative wavefront, the phase speed of the wavefront depends on the depth of the excitation source, with deeper sources producing higher phase speeds. Because the source depth in the solar atmosphere is not known a priori, and constraining it is one of the primary goals of the observations we propose, the cadence must accommodate a plausible range of propagation speeds. For wavefront perturbation speeds between 7 km/s, corresponding to the lower bound of the photospheric sound speed, and 11 km/s, the upper range observed in the simulations analyzed here, sampling intervals of between 1.5 and 0.95 s satisfy the critical sampling requirement. Therefore, we suggest monochromatic sampling at $\lambda^{**}$ with a cadence of 0.5 s, which over-samples the expected wavefront propagation while maintaining a high SNR of approximately 817. Post-processing can further increase the effective SNR through temporal averaging (e.g., a running average).
@@ 6.2 Multi-wavelength Observing Strategy
The wavelength of maximum sensitivity $\lambda^{*}(t)$ in the blue shoulder of Fe I 6302.5 (in Figure 5) extends from 6302.419 Å and 6302.465 Å as the wavefront propagates upward in the atmosphere over a time period of 30 s, with a slope of $1.621\times 10^{-03}\,$Å/s. The VTF IPC imposes a minimum wavelength step between successive observations of 0.0315 Å. Observations sampling $\lambda^{*}(t)$ would ideally be spaced closer in wavelength, but under this restriction, the strategy is to interleave observations over three wavelengths, $\lambda^{*}(t)$ and $\lambda^{**}\pm 0.0315$. The wavelengths to observe are then 6302.3935, 6302.425,and 6302.4565 Å. This can be done with a cadence of 0.48 s achieving a signal to noise at each wavelength of 408. Since the endpoints of this wavelength triplet lie slightly outside of the $\lambda^{*}(t)$ range, and thus have lower wavefront sensitivity, an alternative two wavelength option is warranted, observing the lower end of the $\lambda^{*}(t)$ range at 6302.42 Å and 6302.45 Å. This two-wavelength observation can be done with a cadence of 0.48s with a signal to noise of 527. The propagation of the wavefront between these two wavelengths takes about 18.5s.
An observation period of 60 s with excellent seeing would allow approximately two full wave periods to be sampled under the first multiwavelength scheme and approximately three under the second. Longer excellent seeing time series would be even more beneficial. In the MURaM photosphere, we observe about 50 propagating wavefront events per hour over the $36\,{\rm Mm}^{2}$ field. Over the full $60^{\prime\prime}\times 60^{\prime\prime}$ VTF field of view we thus expect to observe about 40 events per minute.
As discussed above, post observational analysis of the multi wavelength data cube is critical and should focus on using the height-time dependencies of wavefront trajectories to increase detection likelihood. This can be done by looking directly for propagating signals in the four dimensional space-wavelength-time data cube or by shifting the data accumulated at different wavelengths and times so that they can be added together to reinforce the signal. Since the data is to be acquired at a cadence of 0.480.48\,s and the propagation time across the sampled wavelengths takes about 15 to 30 s for the two observation scenerios,trajectories with different upward signal propagation speeds should be accessible and distinguishable in the data cube.
@@ 7 Conclusions
In this work, we have developed a physically motivated framework for detecting locally excited wavefronts in the solar photosphere by connecting the source wavefronts to their wavelength-dependent observational signatures in commonly observed Fe I spectral lines. This connection was established by determining the wavelengths of maximum sensitivity to the wavefront perturbation as it propagates upward into the overshoot region of a simulated solar atmosphere. Based on these synthetic results we defined two observation strategies for the detection of these transient events.
While the detailed results derived are based on a single representative acoustic source, our preliminary examination on more than 25 acoustic events indicates that the qualitative behavior reported here, particularly the speed and extent of wavefront propagation, is common across many of these sources. A comprehensive ensemble analysis spanning the full source population is currently in progress and will be presented in future work.
@@
A central conclusion of the work is that the measurement techniques we sugggest are well fit to DKIST/VTF capabilities, though a smaller minimum wavelength separation between successive observations would be very beneficial. While the detection of expanding, spatially coherent wavefronts is incompatible with raster-based spectroscopy, which entangles spatial structure with temporal evolution and thus confuses signals from fast transient events, the VTF’s Fabry–Pérot–based, full-field imaging spectroscopy preserves their coherence. Moreover the temporal offset between time-series of interleaved wavelengths, can be leveraged to enhance the propagating signal. The VTF performance calculations suggest that, under the VTF operational constraints, sparse wavelength sampling concentrated in the most sensitive line-wing regions can satisfy the temporal requirements imposed by signal propagation, while still retaining sufficient spectral information to track the upward evolution implied by the drift of the optimal sensitivity wavelength.
@@
At the same time, the approach carries inherent limitations that arise directly from its design. First, the usability of the wavefront identification technique depends sensitively on the stability and alignment of successive image frames. Wavefront identification is based on detecting coherent, rapidly evolving perturbations against the more slowly varying background, such that small residual differences between consecutive frames can produce a difference signal. Although post-image reconstruction can help, it cannot remove image misalignment and instability entirely and therefore these effects impose practical limits on the minimum detectable wave amplitude.
@@
Second, the requirement for sub-second to second temporal cadence, necessary to robustly uncover the wavefront signal from the noise by temporal differencing, constrains the number of wavelength positions that can be sampled within a single scan. Consequently, the method trades spectral coverage for temporal sampling. This, combined with practical restrictions on the minimum wavelength separation between successive observations, limits the ability to perform detailed atmospheric inversions or to disentangle overlapping contributions from different heights.
@@
Third, spectral resolution and stability play a central role in determining the vertical localization of the measurement. Our current method relies on isolating wave-induced perturbations over a narrow and stable wavelength range, as encoded in the wavelength-dependent response functions of the spectral line. If the instrument resolution or stability worsens, contributions from a wider range of atmospheric heights are mixed within a single wavelength sample, reducing the contrast of the wavefront signal relative to background fluctuations. In this regard, spectral lines with soft (not steep) shoulder more reliably localize measurements in height.
@@
Despite these limitations, our assessment indicates that the DKIST 4 m aperture, in combination with the VTF Fabry–Pérot imager, constitutes one of the few observational platforms capable of simultaneously satisfying the spatial, spectral, and temporal resolution constraints required by the method developed here. Consequently, local acoustic wavefront detection represents both a compelling scientific objective and a stringent benchmark for DKIST’s ability to probe weak, rapidly evolving photospheric dynamics at their intrinsic spatial and temporal scales. Successful application of the framework developed here would transform a long-standing problem into an observationally accessible diagnostic.
@@ Acknowledgments
Sincere thanks to M. Rempel for his continued and ongoing scientific generosity. The simulation material is based upon work supported by the NSF National Center for Atmospheric Research, which is a major facility sponsored by the U.S. National Science Foundation under Cooperative Agreement No. 1852977.

### v2

@@ 6 Observational Strategies – Implications for DKIST/VTF
The two measures, $\lambda^{*}(t)$ and $\lambda^{**}$, motivate different observing strategies. The first is an estimate for the best wavelength sequence to employ to follow the wavefront propagation, while the second is an estimate of the best overall choice of wavelength at which to perform monochromatic imaging at high cadence. Observing at $\lambda^{**}$ is roughly equivalent to observing the wavefronts as they propagate horizontally at the height of maximum sensitivity. Observing sequentially over the wavelengths $\lambda^{*}(t)$, on the other hand, allows collection of a data cube that can be post-processed to find wavefront signatures as they propagate upward with height. For a more robust measurement, the individual $\lambda^{*}$ maps in multi-wavelength observations could be shifted in time, stacked, and summed to obtain the mean signal over all heights. In this operation, the wave signal adds coherently, while the variation in granulation with height contributes to some cancellation of the background and consequent enhancement of the propagating wavefront signal-to-noise.
@@
The VTF instrument provides full-field imaging spectroscopy across 520087005200–8700  Å, with a $60^{\prime\prime}\times 60^{\prime\prime}$ field of view sampled at $0.014^{\prime\prime}$ per pixel, corresponding to about $10.5$ km on the solar surface. This sampling is comparable to the 16 km horizontal grid spacing of the MURaM simulation analyzed for this paper, and allows access to the spatial scales over which acoustic wavefronts remain coherent and may be recoverable by temporal differencing (Bahauddin and Rast, 2023). The spectral resolving power of VTF is $R\sim 10^{5}$ (e.g., $\sim$0.06 Å resolution with $\sim$0.03 Å sampling near 6000 Å) over a free spectral range of $\sim$10 Å. In detail, the VTF spectral sampling width is about 0.023 Å at 5250 Å and 0.033 Å at 6302 Å (Tritschler et al., 2016; Schmidt et al., 2016). These band passes are over-plotted in the insets of Figure 7.
@@
As discussed above, the distributions of $\lambda^{**}$, as determined from the individual atmospheres within the annulus of interest, have widths of $\sigma_{\lambda^{**}}=0.0031$ Å, $\sigma_{\lambda^{**}}=0.0028$ Å, and $\sigma_{\lambda^{**}}=0.0046$ Å  for Fe I 5250.2, Fe I 5250.6, and Fe I 6302.5, respectively. Thus, the range of $\lambda^{**}$ values found for each of these lines lies well within one spectral sampling element of the VTF. VTF observations centered on $\lambda^{**}$ or $\lambda^{*}$ will capture a relatively broad range of wavelengths and consequently sample a range of atmospheric heights. It is clear from Figure 7 that, in the case of Fe I 5250.2 for example, centering the VTF passband on $\lambda^{\ast\ast}$ will include significant contributions from the continuum. Those contributions are dominated by granulation noise that will likely overpower the wave signal in monochromatic observations. More generally, the range of atmospheric heights encompassed by the VTF bandpass may smear the wave signal. However, the blue wing of Fe I 5250.6, and even more so that of Fe I 6302.5, are shallower and thus less contaminated by contributions away from the wavelength of maximum sensitivity (see insets of Figure 7). Here we focus our observational plans on the Fe I 6302.5 line, though careful convolution with both the VTF spectral bandpass and the DKIST point spread function (PSF) is required to fully assess the potential of each line for local wavefront discovery.
@@ 6.1 Monochromatic Observations
To comply with the requirements of the temporal difference method, the cadence of the monochromatic observations should be at or below the critical sampling rate $\Delta t=\Delta x/v_{\mathrm{ph}}$, where $\Delta x$ is the spatial sampling width and $v_{\mathrm{ph}}$ is the phase speed of the wavefront. The phase speed of the wavefront at z=0.152z=0.152\,Mm, the midpoint of the Fe I 6302.5 $\lambda^{**}$ height, is 10.5~10.5 km/s. This speed is estimated from the frame-to-frame radial advance of the wavefront measured with the azimuthally averaged acoustic mask (Figure 3) and the mean value over the interval in which the front traverses the granule is quoted. Critical sampling of the wavefront thus occurs at a cadence of 1 s. The practical viability of this cadence is governed by the signal-to-noise requirements based on the intensity fluctuations caused by acoustic perturbations.
@@
Quantitative guidance on the achievable cadence vs. signal-to-noise trade space is assessed by the VTF Instrument Performance Calculator (IPC, National Solar Observatory, 2026). For monochromatic (single-wavelength) intensity imaging, the IPC indicates that an SNR of 235 can be achieved at a cadence of 0.04 s. Relaxing the cadence to 1.04 s, however, increases the SNR to 1200. While our analysis is for a single representative wavefront, the phase speed of the wavefront depends on the depth of the excitation source, with deeper sources producing higher phase speeds. Because the source depth in the solar atmosphere is not known a priori, and constraining it is one of the primary goals of the observations we propose, the cadence must accommodate a plausible range of propagation speeds. For wavefront propagation speeds between 7 km s-1 (the lower bound of the photospheric sound speed) and 11 km s-1 (the upper range observed in the simulations analyzed here), sampling intervals between 0.95 and 1.5 s satisfy the minimum sampling requirement. Therefore, we suggest monochromatic sampling at $\lambda^{**}$ with a cadence of 0.5 s, which over-samples the expected wavefront propagation while maintaining a high SNR of approximately 817. Post-processing can further increase the effective SNR through temporal averaging (e.g., a running average).
@@ 6.2 Multi-wavelength Observing Strategy
The wavelength of maximum sensitivity $\lambda^{*}(t)$ in the blue shoulder of Fe I 6302.5 (Figure 6) extends from 6302.419 Å to 6302.465 Å as the wavefront propagates upward in the atmosphere over a time period of 30 s, with a slope of $1.621\times 10^{-03}\,$Å s-1. The VTF IPC imposes a minimum wavelength step between successive observations of 0.0315 Å. Observations sampling $\lambda^{*}(t)$ would ideally be spaced closer in wavelength but, under this restriction, the strategy is to interleave observations over three wavelengths, $\lambda^{**}$ and $\lambda^{**}\pm 0.0315$, i.e., 6302.3935, 6302.425, and 6302.4565 Å. This can be done with a cadence of 0.48 s achieving a signal to noise at each wavelength of 408. Since the endpoints of this wavelength triplet lie slightly outside of the $\lambda^{*}(t)$ range, and thus have lower wavefront sensitivity, an alternative two wavelength option is warranted, observing the lower end of the $\lambda^{*}(t)$ range at 6302.42 Å and 6302.45 Å. This two-wavelength observation can be done with a cadence of 0.48 s with a signal to noise of 527. The wavefront transit time between the atmospheric heights sampled by these two wavelengths is about 18.5 s.
An observation period of 60 s with excellent seeing would allow approximately two full wavefront transit times to be sampled under the first multi-wavelength scheme and approximately three under the second. Longer excellent seeing time series would further improve the sampling statistics. In the MURaM photosphere, we observe about 25 propagating wavefront events per hour over the $36\,{\rm Mm}^{2}$ field. Thus, we expect to observe about 40 events per minute over the full $60^{\prime\prime}\times 60^{\prime\prime}$ VTF field of view (FOV). This estimate is an upper limit that depends on the detection criteria and assumes maximum possible spatial resolution over the whole FOV, which is not guaranteed. However, even with an order of magnitude fewer events, we are still poised for multiple detections per minute.
As discussed above, post-observational analysis of the multi-wavelength data cube is critical and should focus on using the height-time dependencies of wavefront trajectories to increase detection likelihood. This can be done by looking directly for propagating signals in the four-dimensional space-wavelength-time data cube or by shifting the data accumulated at different wavelengths and times so that they can be added together to reinforce the signal. Since the data is to be acquired at a cadence of 0.480.48\,s and the propagation time across the sampled wavelengths takes about 15 to 30 s for the two observation scenerios, trajectories with different upward signal propagation speeds should be accessible and distinguishable in the data cube.
@@ 7 Conclusions
In this work, we have developed a physically motivated framework for detecting locally excited wavefronts in the solar photosphere by connecting the source wavefronts to their wavelength-dependent observational signatures in commonly observed Fe I spectral lines. This connection was established by determining the wavelengths of maximum sensitivity to the wavefront perturbation as it propagates upward through the overshoot region of a simulated solar atmosphere. Based on these synthetic results, we defined two observation strategies for the detection of these transient events.
@@
A central conclusion of the work is that the measurement techniques we suggest are well fit to DKIST/VTF capabilities, though a smaller minimum wavelength separation between successive observations would be very beneficial. While the detection of expanding, spatially coherent wavefronts is incompatible with raster-based spectroscopy, which entangles spatial structure with temporal evolution and thus confuses signals from fast transient events, the VTF’s Fabry–Pérot–based, full-field imaging spectroscopy preserves their coherence. Moreover, the temporal offset between time series of interleaved wavelengths can be leveraged to enhance the propagating signal. VTF performance calculations suggest that, under the operational constraints of the instrument, sparse wavelength sampling concentrated in the most sensitive line-wing regions can satisfy the temporal requirements imposed by signal propagation while still retaining sufficient spectral information to track the upward evolution implied by the drift of the optimal sensitivity wavelength.
@@
At the same time, the approach carries inherent limitations that arise directly from its design. First, the usability of the wavefront identification technique depends sensitively on the stability and alignment of successive image frames. Wavefront identification is based on detecting coherent, rapidly evolving perturbations against the more slowly varying background, such that small residual differences between consecutive frames can produce a difference signal. Although post-image reconstruction (e.g., Van Noort et al., 2005) can help, it cannot remove image misalignment and instability entirely. Therefore, these effects impose practical limits on the minimum detectable wave amplitude. A comprehensive test of the whole data acquisition process, including simulating time-dependent atmospheric seeing and an end-to-end test of the image reconstruction for VTF-like data, would help assess the impact of image stability on wavefront detection, but such a test goes beyond the scope of this paper.
@@
Second, the critical requirement for sub-second to second temporal cadence to robustly separate the wavefront signal from the noise by temporal differencing constrains the number of wavelength positions that can be sampled within a single scan. Consequently, the method trades spectral coverage for temporal sampling. Combined with practical restrictions on the minimum wavelength separation between successive observations, this limits the ability to perform detailed atmospheric inversions or to disentangle overlapping contributions from different heights.
@@
Third, spectral resolution and stability play a central role in determining the vertical localization of the measurement. Our current method relies on isolating wave-induced perturbations over a narrow and stable wavelength range, as encoded in the wavelength-dependent response functions of the spectral line. If the instrument resolution or stability worsens, contributions from a wider range of atmospheric heights are mixed within a single wavelength sample, reducing the contrast of the wavefront signal relative to background fluctuations. In this regard, spectral lines with soft (not steep) wings more reliably localize measurements in height. Our preliminary tests show that the Na I D lines at 5888 Å and 5896 Å exhibit such features, thanks to their broad wings as compared to the Fe spectral lines considered in this paper. Technical observations in one of the Na lines were already conducted as a part of the VTF’s first light (Schubert et al., 2025). Given their non-LTE nature, the Na I D lines require more detailed modeling, which is currently in progress.
@@
Despite these limitations, our assessment indicates that the DKIST 4 m aperture, in combination with the VTF Fabry–Pérot imager, constitutes one of the few observational platforms capable of simultaneously satisfying the spatial, spectral, and temporal resolution constraints required by the methods suggested here. Consequently, local acoustic wavefront detection represents both a compelling scientific objective and a stringent benchmark for DKIST’s ability to probe weak, rapidly evolving photospheric dynamics at their intrinsic spatial and temporal scales. Successful application of the framework developed would transform a long-standing problem into an observationally accessible diagnostic.
@@ Acknowledgments
Sincere thanks to M. Rempel for his continued and ongoing scientific generosity. The simulation material is based upon work supported by the NSF National Center for Atmospheric Research, which is a major facility sponsored by the U.S. National Science Foundation under Cooperative Agreement No. 1852977. The authors thank the anonymous referees for their careful reading and valuable input.